# Exploring the Potential of Encoder-free Architectures in 3D LMMs

### v2

@@ 4 Results and Visualization
Results. In Table 5, on the Objaverse benchmark (Deitke et al., 2023), Enel-7B achieves a GPT score of 50.98% for 3D object captioning, setting a new SOTA. In traditional metrics, Sentence-BERT and SimCSE reach 48.65% and 49.36%, respectively, comparable to ShapeLLM-13B. For 3D object classification, Enel-7B outperformes prior encoder-based 3D LMMs with a GPT score of 55.10%. Given the same training dataset as PointLLM, these results validate the effectiveness of our proposed LLM-embedded Semantic Encoding and Hierarchical Geometry Aggregation strategies for the encoder-free architecture. Additionally, on the 3D-VQA task of the 3D MM-Vet dataset (Qi et al., 2024), despite the lack of spatial and embodied interaction-related data in the training set, Enel achieves the GPT score of 43.1%, surpassing PointLLM-7B by 1.9%. Replacing 7B Vicuna with 13B Vicuna, Enel-13B achieves substantial performance gains across tasks. Details of the evaluation metric is provided in Appendix B.2.
Visualization. In the Figure 5, we visualize the attention scores between the average text token and the point tokens in the last layer of both PointLLM and Enel. Three object categories, including the chair, the airplane, and the desk lamp, are selected from the Objaverse dataset (Deitke et al., 2023). In the Figure 5, red indicates higher values. We observe that in encoder-based 3D LMMs, the semantic relevance between the text tokens and the processed point tokens is relatively low. In contrast, Enel, with its encoder-free architecture, achieves a high correlation between the features of the two different modalities, with the average text token focusing on key geometric structures of the objects, such as the backrest of the chair, the wings of the airplane, and the lampshade of the desk lamp.
@@ 5 Conclusion
Appendix A Experimental Settings
@@
Implementation Details. We use the LLaMA model (Touvron et al., 2023) as our LLM backbone, with the 7B Vicuna-v1.1 (Chiang et al., 2023) checkpoint as the default setting. In the token embedding layer, the point cloud is first processed by a linear layer to expand its dimension from 6 to 288. The input point cloud initially consists of 8192 points, followed by three iterations of farthest point sampling (FPS), reducing the size to 512, 256, and 128, respectively. After each FPS operation, k-Nearest Neighbors (k-NN) is applied with a cluster size of 81. And geometric features are extracted using triangular encoding, followed by linear layers that progressively increase the dimension to 576, 1152, and 2304. Finally, the projection layer maps the features to the LLM dimension of 4096. In the pre-training stage, we unfreeze the first four LLM layers. Within the LLM-embedded Semantic Encoding strategy, Hybrid Semantic Loss applies masked modeling to 30% of the tokens and reconstructs the patches for the remaining 70% visible tokens. During instruction tuning, geometric aggregation is applied at the end of the 1st, 2nd, and 3rd LLM layers, reducing point tokens from 128 to 64, 32, and 16. MaxMean pooling is used to retain more information. After two LLM layers, geometric propagation is applied at the end of the 6th, 7th, and 8th layers to restore the number of point cloud to 128.
@@
Training and Evaluation Details. During the two-stage training, each stage utilizes the same dataset and preprocessing method as PointLLM. All training are conducted on 4 $\times$ 80G A100 GPUs in BF16 precision, utilizing FlashAttention, the AdamW optimizer, and a cosine learning rate schedule. During the pre-training stage, the model is trained for three epochs with a batch size of 128 and a learning rate of 4e-4. In the instruction tuning stage, it is conducted for three epochs with batch size of 32 and a learning rate of 2e-5. The GPT-4 model (Achiam et al., 2023) used for classification and captioning tasks evaluation refers to “gpt-4-0613” version consistent with PointLLM (Xu et al., 2023). In contrast, the GPT-4 model employed for QA performance evaluation corresponds to “gpt-4-0125” version aligning with ShapeLLM (Qi et al., 2024). Additionally, the GPT evaluation prompts for classification and captioning are identical to those used in PointLLM, while the prompts for QA follow those in ShapeLLM.
@@
Table 6: Ablation Experiments. We begin the ablation experiments by changing the single configuration of the module from ENEL. Psi and Phi denote mask ratios of 60% and 30%, respectively. For Hybrid Semantic Loss, the subscripts patch and feat refer to the masked modeling target, with reconstruction targeting the corresponding feat and patch. $l$ represents the number of aggregation and propagation operations. $H$ refers to the LLM layers between $l$ aggregation and $l$ propagation operations. $O$ refers to the LLM layer between two individual aggregation or propagation operations.
Model Cap Cls (Avg)
@@
ENEL-7B 50.98 48.65 49.36 3.85 7.20 12.60 55.10
–Hybrid Semantic Loss 47.15 48.06 48.31 3.40 7.43 11.84 50.50
Hybrid Semantic LosspatchΦ 49.13 48.80 49.20 3.99 7.20 12.38 52.30
Hybrid Semantic LosspatchΨ 48.79 48.30 49.00 3.65 6.90 11.98 52.10
Hybrid Semantic LossfeatΨ 49.62 48.00 48.67 3.78 6.82 12.33 51.50
–gate mechanism 49.61 48.41 48.97 3.79 7.12 12.48 53.60
$l$=2, $H$=2, $O$=0 48.83 48.20 48.53 3.72 6.89 12.01 51.50
$l$=2, $H$=4, $O$=0 49.05 48.47 48.62 3.65 7.10 12.31 52.20
$l$=2, $H$=2, $O$=2 48.96 47.95 48.88 3.80 7.05 12.55 52.00
$l$=2, $H$=4, $O$=2 49.68 48.70 48.85 3.84 7.56 12.76 53.10
Appendix B More Experiments
B.1 Variants of Point Cloud Self-Supervised Learning Losses.
@@ Appendix B More Experiments
As seen in Figure 7 (a), in the Masked Modeling Loss, after the learnable tokens are processed by the LLM, the tokens are transformed to the point patches $\{G_{\text{pre}_{i}}\}_{i=1}^{M*r}\in\mathbb{R}^{M*r\times k\times 3}$ through a linear layer. We utilize the $l_{2}$ chamfer distance to align the predicted $G_{\text{pre}}$ with the point patches $G_{mask}$ corresponding to the masked tokens, reconstructing the spatial information. The optimization can be written as
@@
Finally, in the Figure 7 (c) Hybrid Semantic Loss, the masked tokens and the corresponding patches are referred to as $\{F_{\text{mask}_{i}}\}_{i=1}^{M*r}$ and $\{G_{\text{mask}_{i}}\}_{i=1}^{M*r}$, respectively. The remaining tokens are denoted as$\{F_{\text{vis}_{i}}\}_{i=1}^{M*(1-r)}$ and $\{G_{\text{vis}_{i}}\}_{i=1}^{M*(1-r)}$. After passing point tokens through the LLM, we compute the MSE between $F_{\text{pre}}$ and $F_{\text{vis}}$. The learnable tokens $F_{\text{learn}}$ are transformed into $G_{\text{pred}}$, and the $L_{2}$ Chamfer distance is computed between $G_{\text{pred}}$ and $G_{\text{mask}}$. These two are added to the original cross-entropy loss with coefficients all equal to 1.
B.2 Metric Analysis
@@
B.3 More Ablation Experiments
@@
The Effects of LLM-embedded Semantic Encoding Strategy. In the Table 6, on the basis of Enel, removing the Hybrid Semantic Loss during the pre-training stage significantly degrades performance. The GPT-4 score for the captioning task drops from 50.98% to 47.15%, and the GPT-4 score for the classification task decreases to 50.50%. This is because the proposed self-supervised learning function for point clouds effectively captures the detailed structures and high-level semantics.
@@
Appendix C Model Output

### v4

@@ 4 Results and Visualization
Results. In Table 5, on the Objaverse benchmark Deitke et al. (2023), Enel-7B achieves a GPT score of 51.03% for 3D object captioning, setting a new SOTA. In traditional metrics, Sentence-BERT and SimCSE reach 48.79% and 49.52%, respectively, comparable to PointLLM-PiSA-13B. For 3D object classification, Enel-7B outperformes prior encoder-based 3D LMMs with a GPT score of 55.55%. Given the same training dataset as PointLLM, these results validate the effectiveness of our proposed LLM-embedded Semantic Encoding and Hierarchical Geometry Aggregation strategies for the encoder-free architecture. Additionally, on the 3D-VQA task of the 3D MM-Vet dataset Qi et al. (2024), despite the lack of spatial and embodied interaction-related data in the training set, Enel achieves the GPT score of 43.8%, surpassing PointLLM-7B by 2.6%. Replacing 7B Vicuna with 13B Vicuna, Enel-13B achieves substantial performance gains across tasks. When replacing the Vicuna-7B with Qwen2.5-7B and using ShapeLLM training data, Enel-7B achieves over 6% improvements across benchmarks. Details of the evaluation metric and classification performance on the ModelNet dataset are provided in Appendix A.3.2 and A.3.3, respectively.
Visualization. In the Figure 5, we visualize the attention scores between the average text token and the point tokens in the last layer of both PointLLM and Enel. Three object categories, including the chair, the airplane, and the desk lamp, are selected from the Objaverse dataset Deitke et al. (2023). In the Figure 5, red indicates higher values. We observe that in encoder-based 3D LMMs, the semantic relevance between the text tokens and the processed point tokens is relatively low. In contrast, Enel, with its encoder-free architecture, achieves a high correlation between the features of the two different modalities, with the average text token focusing on key geometric structures of the objects, such as the backrest of the chair, the wings of the airplane, and the lampshade of the desk lamp.
@@ 5 Conclusion
Appendix A Appendix
A.1 Related Work
3D LMM. Recent advancements in integrating large language models (LLMs) with 3D data have led to significant progress in both object-level and scene-level understanding. At the object level, early approaches like Hong et al. (2024) utilize 2D rendering to leverage 2D LLMs, but this sacrifices geometric details. More recent models, including Point-Bind LLM Guo et al. (2023), PointLLM Xu et al. (2023b) and ShapeLLM Qi et al. (2024), directly encode point clouds and align them with LLMs, by combining the 3D encoder with a powerful language model, effectively fusing geometric, appearance, and linguistic information. At the scene level, models like Chat-3D Wang et al. (2023) and Scene-LLM Fu et al. (2024) focus on understanding complex spatial relationships through dialogue and tasks like captioning. Scene-LLM Fu et al. (2024) enhances embodied agents’ abilities in interactive 3D indoor environments by integrating both scene-level and egocentric 3D information. Grounded 3D-LLM Chen et al. (2024b) utilizes referent tokens to reference specific objects within 3D scenes, enabling tasks such as object detection and language grounding.
Encoder-free Vision-Language Models. Traditional vision-language models (VLMs) often rely on vision encoders to extract visual features before processing them with language models, integrating image encoders like CLIP Radford et al. (2021) and DINO V2 Oquab et al. (2023). However, recent efforts have explored encoder-free VLMs for their simplicity. Approaches like ChameleonTeam (2024); Xie et al. (2024) use VQ tokenizers Esser et al. (2021) or linear projection layers Diao et al. (2024a); Chen et al. (2024a) to represent images. Fuyu-8B Bavishi et al. (2023), a pure decoder-only model, directly processes image patches through linear projections, handling high-resolution images but showing only average performance. The EVE series Diao et al. (2024b; 2025) eliminates the need for a separate vision encoder by bridging vision-language representation within a unified decoder and enhancing visual recognition capabilities through additional supervision. Mono-InternVL series Luo et al. (2024; 2025) leverage visual experts and progressive visual pre-training (EViP/EViP++) to achieve stable optimization and competitive performance. SAIL series Lei et al. (2025) directly encode raw pixels and decodes language within a single architecture, achieving competitive vision-language performance without pre-trained vision encoders.
Figure 7: Variants of Point Cloud Self-Supervised Learning Losses. (a) The Variant of Masked Modeling Loss, (b) The Variant of Reconstruction Loss, (c) The Variant of Hybrid Semantic Loss.
A.2 Experimental Settings
@@
Implementation Details. We use the LLaMA model Touvron et al. (2023) as our LLM backbone, with the 7B Vicuna-v1.1 Chiang et al. (2023) checkpoint as the default setting. In the token embedding layer, the point cloud is first processed by a linear layer to expand its dimension from 6 to 288. The input point cloud initially consists of 8192 points, followed by three iterations of farthest point sampling (FPS), reducing the size to 512, 256, and 128, respectively. After each FPS operation, k-Nearest Neighbors (k-NN) is applied with a cluster size of 81. And geometric features are extracted using triangular encoding, followed by linear layers that progressively increase the dimension to 576, 1152, and 2304. Finally, the projection layer maps the features to the LLM dimension of 4096. In the pre-training stage, we unfreeze the first four LLM layers. Within the LLM-embedded Semantic Encoding strategy, Hybrid Semantic Loss applies masked modeling to 30% of the tokens and reconstructs the patches for the remaining 70% visible tokens. During instruction tuning, geometric aggregation is applied at the end of the 1st, 2nd, and 3rd LLM layers to reduce point tokens. MaxMean pooling is used to retain more information. After two LLM layers, geometric propagation is applied at the end of the 6th, 7th, and 8th layers to restore the number of point cloud to 128. After two LLM layers, geometric aggregation is applied at the 11th–13th layers, followed by geometric propagation at the 16th–18th layers.
@@
Training and Evaluation Details. During the two-stage training, each stage utilizes the same dataset and preprocessing method as PointLLM. All training are conducted on 4 $\times$ 80G A100 GPUs in BF16 precision, utilizing FlashAttention, the AdamW optimizer, and a cosine learning rate schedule. During the pre-training stage, the model is trained for three epochs with a batch size of 128 and a learning rate of 4e-4. In the instruction tuning stage, it is conducted for three epochs with batch size of 32 and a learning rate of 2e-5. The GPT-4 model Achiam et al. (2023) used for classification and captioning tasks evaluation refers to “gpt-4-0613” version consistent with PointLLM Xu et al. (2023b). In contrast, the GPT-4 model employed for QA performance evaluation corresponds to “gpt-4-0125” version aligning with ShapeLLM Qi et al. (2024). Additionally, the GPT evaluation prompts for classification and captioning are identical to those used in PointLLM, while the prompts for QA follow those in ShapeLLM.
@@
Table 6: Ablation Experiments. We begin the ablation experiments by changing the single configuration of the module from Enel. $\Psi$ represents a mask ratio of 60%, while $\Phi$ represents a mask ratio of 30%. For Hybrid Semantic Loss, the subscript patchpatch and featfeat represent the masked modeling target, while the reconstruction target is the corresponding featfeat and patchpatch. $l$ represents the number of aggregation and propagation operations. $H$ refers to the LLM layers between $l$ aggregation and $l$ propagation operations. $O$ refers to the LLM layer between two individual aggregation or propagation operations.
Model Cap Cls
@@
Enel-7B 50.92 48.61 49.31 3.88 7.20 12.50 55.00
– Hybrid Semantic Loss 47.19 48.07 48.31 3.46 7.41 11.84 50.61
Hybrid Semantic LosspatchΦ 49.05 48.82 49.20 4.01 7.25 12.38 52.20
Hybrid Semantic LosspatchΨ 48.96 48.38 49.00 3.66 6.97 11.98 52.00
Hybrid Semantic LossfeatΨ 49.63 48.00 48.62 3.78 6.88 12.33 51.50
– gate mechanism 49.26 48.41 48.93 3.71 7.12 12.47 53.50
l=2,H=2,O=0 48.81 48.10 48.57 3.70 6.99 12.01 51.50
l=2,H=4,O=0 49.02 48.47 48.61 3.65 7.10 12.31 52.00
l=2,H=2,O=2 48.96 47.96 48.89 3.80 7.05 12.55 52.00
l=2,H=4,O=2 49.58 48.70 48.84 3.84 7.56 12.76 53.00
A.3 More Experiments
A.3.1 Variants of Point Cloud Self-Supervised Learning Losses.
@@ A.3 More Experiments
As seen in Figure 7 (a), in the Masked Modeling Loss, after the learnable tokens are processed by the LLM, the tokens are transformed to the point patches $\{G_{\text{pre}_{i}}\}_{i=1}^{M*r}\in\mathbb{R}^{M*r\times k\times 3}$ through a linear layer. We utilize the $l_{2}$ chamfer distance to align the predicted $G_{\text{pre}}$ with the point patches $G_{mask}$ corresponding to the masked tokens, reconstructing the spatial information. The optimization is:
@@
[…]
Table 7: Comparison of computational complexity between PointLLM-7B and ENEL-7B. S1 and S2 refer to the pre-training and instruction tuning stages, respectively. Conv. Steps indicates the number of steps required for loss convergence.
Method Time (H) Memory (S1/S2) FLOPs Conv. Steps (S1/S2)
PointLLM-7B 31.6 67G / 57G $2.0\times 10^{18}$ 10100 / 4300
ENEL-7B 22.2 56G / 42G $1.59\times 10^{18}$ 9790 / 3700
Improvement 29.7% 16.4% / 26.3% 20.5% 2.9% / 14.0%
A.3.2 Metric Analysis
@@
Table 8: ModelNet40 classification results under instruction-typed and completion-typed prompts. The instruction-typed (I) prompt is “What is this?” and the completion-typed (C) prompt is “This is an object of.”
Model ModelNet (I) ModelNet (C) ModelNet-Avg
PointLLM-7B 53.44 51.82 52.63
PointLLM-13B 53.00 52.55 52.78
ShapeLLM-7B 53.08
ShapeLLM-13B 52.96
PointLLM-PiSA-7B 54.58 52.60 53.59
PointLLM-PiSA-13B 55.03 53.81 54.42
Enel-7B 54.82 53.69 54.26
Enel-13B 55.59 54.38 55.00
Enel-7B* 61.25 60.47 60.86
A.3.3 ModelNet Classification Task
As shown in Table 8, Enel-7B achieves an average accuracy of 54.26%, surpassing PointLLM-7B (52.63%), ShapeLLM-7B (53.08%) and PointLLM-PiSA-7B (53.59%). Similarly, ENEL-13B reaches 55.00%, outperforming both ShapeLLM-13B (52.96%) and PointLLM-PiSA-13B (54.42%). These results demonstrate the effectiveness of the encoder-free design in 3D object understanding.
A.3.4 Complexity Analysis
In Table 7, compared to PointLLM-7B, Enel-7B demonstrates significant improvements while using the same training dataset. It achieves 29.7% faster training time, reduces GPU memory usage by 16.4% and 26.3% in Stage 1 and Stage 2, respectively, lowers training FLOPs by 20.5%, and accelerates convergence speed by 2.9% (Stage 1) and 14.0% (Stage 2).
A.3.5 Encoder-Free Architecture Claim.
Following the consensus in recent Large Multimodal Model (LMM) literature, we strictly define an architecture as “encoder-free” based on two criteria: (1) the absence of a heavy, independently pretrained visual backbone, and (2) the utilization of end-to-end training from scratch. Unlike traditional 3D LMMs that rely on decoupled, pretrained encoders (e.g., Point-BERT Yu et al. (2022)) for semantic extraction, our design integrates a lightweight, randomly initialized embedding layer trained jointly with the LLM.
Alignment with Community Standards. This design philosophy parallels established encoder-free paradigms in the 2D image and video domains. For instance, EVE Diao et al. (2024a) utilizes a token embedding layer based on convolution and cross-attention ($\sim$16M parameters), while ELVA Li et al. (2025) employs a spatio-temporal attention layer ($\sim$9M parameters) for video framing. Similarly, Mono-InternVL Luo et al. (2024) relies on a lightweight stack of convolutions ($\sim$10M parameters). As detailed in Table 9, our proposed point embedding layer comprises only 3M parameters. This is not only significantly more lightweight than its 2D counterparts but also orders of magnitude smaller than typical 3D encoders (e.g., $\sim$88M for PointBERT used in PointLLM). Our module functions strictly as a tokenizer rather than a visual encoder.
Table 9: Comparison of Tokenizer Parameters across Domains.
Method Domain Tokenizer Structure Tokenizer Params Ratio (Tok./Total)
EVE / EVEv2 Diao et al. (2024a; 2025) Image Conv + Cross-Attn 16 M $\sim$0.23%
Mono-InternVL Luo et al. (2024) Image Stacked Conv 10 M $\sim$0.14%
ELVA Li et al. (2025) Video Spatio-temporal Attn 9 M $\sim$0.13%
PointLLM Xu et al. (2023a) 3D PointBERT Encoder $\sim$88 M $\sim$1.24%
Ours 3D Point Embedding 3 M $\sim$0.04%
Structural Formatting vs. Semantic Encoding. We explicitly distinguish the structural operations used in our embedding layer—specifically Farthest Point Sampling (FPS) and $k$-Nearest Neighbors ($k$-NN)—from semantic encoding. Due to the data irregularity of unstructured 3D point clouds, FPS and $k$-NN serve as the mathematically necessary equivalents of the “patchify” or “stride” operations used in 2D Vision Transformers. They are required to group raw data points into processable tokens. Crucially, these operations are parameter-free. The subsequent learnable MLPs serve only to project these local geometric groupings into the feature dimension required by the LLM.
A.3.6 More Ablation Experiments
@@
The Effects of LLM-embedded Semantic Encoding Strategy. In the Table 6, on the basis of Enel, removing the Hybrid Semantic Loss during the pre-training stage significantly degrades performance. The GPT-4 score for the captioning task drops from 51.03% to 47.15%, and the GPT-4 score for the classification task decreases to 50.50%. This is because the proposed self-supervised learning function for point clouds effectively captures the detailed structures and high-level semantics.
@@
A.4 Use of LLM
In this paper, the Large Language Model (LLM) is mainly used for grammar checking and spelling error detection.
A.5 Model Output